OpenStudy (anonymous):

Find the slope intercept form of the line whose slope is 4 and that passes through the point (-7,8) the equation of the line is?

OpenStudy (mrnood):

you know that the slope - intercept equation looks like this: y=mx+b OK?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

That's as far as I understand... I don't know how to get the answer.

OpenStudy (mrnood):

ok well the 'm' in that equation is the slope the question tells oyu the slope so can you start to write the equation - putting th ecorrect value for m (but leave 'b' for now...)

OpenStudy (anonymous):

Y=4x

OpenStudy (mrnood):

y = mx +b m = 4 so write the equation with the value for m in it

OpenStudy (mrnood):

(you need the "+b" on the end )

OpenStudy (anonymous):

Ok. How do you get b?

OpenStudy (anonymous):

Oh, y=4x+b

OpenStudy (mrnood):

at the moment I just wanted you to write the equation like this y = 4x+b we still need to find b OK? - we are doing this in 2 steps - you have done one of them so far Do oyu understand why the equation starts to be y=4x+b ?

OpenStudy (anonymous):

Yes..

OpenStudy (mrnood):

OK good So now the second step is to use the second bit of information from the question that is the point (-7,8) lise on the line this means that the equation is true when x=-7, and y = 8 so write the equation y = 4x+b with the values x= -7 and y = 8 in it (leave b as it is for now)

OpenStudy (anonymous):

Okay..

OpenStudy (mrnood):

so write the equation y = 4x+b x= -7 y = 8 in it (leave b as it is for now)

OpenStudy (anonymous):

Y=4x+b y =8x=-7 this doesn't make sense to me..

OpenStudy (mrnood):

wherever there is y in the equation write 8 wherever there is x write -7 (BUT remember that 4x means 4 times x)

OpenStudy (anonymous):

8=4-7+b yes?

OpenStudy (anonymous):

B=5

OpenStudy (mrnood):

yes - but to be clear about 4 times x it is clearer if oyu write 8 = 4(-7) +b

OpenStudy (anonymous):

Y=4(-7)+5. Yes?

OpenStudy (mrnood):

yes so that can be rewritten as 8= -28 +b so what is b?

OpenStudy (anonymous):

Okay.. So I don't need it to equal 5?

OpenStudy (anonymous):

B is 5 no??

OpenStudy (mrnood):

no it's not , please don't rush - I am trying to guide you to work out the answer

OpenStudy (anonymous):

Okay I see..

OpenStudy (anonymous):

Sorry..

OpenStudy (mrnood):

To review what we did.. we KNOW the equation will be in th eform y = mx + b and we know that m is th eslope and we know the slope is 4 so we know th eequaiton is y= 4x+b BUT we also know that the point (-7,8) is on the line (which means the equationis true when it passes through that point) SO we know that when x = -7 then y = 8 so back to the equation we put y - 8 and x = -7 so 8=4×(−7)+b this is the same as 8 = -28 + b so you can work out b

OpenStudy (anonymous):

Yes... B =36

OpenStudy (mrnood):

well done so now write the whole equation y=mx+b with the values for m and b that you calculated

OpenStudy (anonymous):

Y=4x+36

OpenStudy (anonymous):

Correct ?

OpenStudy (mrnood):

correct well done this is the principle of all these questions - once you know the slope, you can use 1 point on the line to find b

OpenStudy (anonymous):

Thank you sir

OpenStudy (anonymous):

So, I attempted my next one and the slope is 7 the point is (-4,11). I came up with y=7x+35 but I was wrong it is 39 and I don't know where I went wrong??

OpenStudy (mrnood):

ok you made a good attempt starting with y = 7x + b then we put in the point values 11= 7 (-4) +b 11 = -28 +b you can see that 39 is the correct value